# Stabilizing Thompson Sampling with Point Null Bayesian Response-Adaptive Randomization

### v1

@@ Abstract
Response-adaptive randomization (RAR) methods use accumulated data to adapt randomization probabilities, aiming to increase the probability of allocating patients to effective treatments. A popular RAR method is Thompson sampling, which randomizes patients proportionally to the Bayesian posterior probability that each treatment is the most effective. However, its high variability early in a trial can also increase the risk of assigning patients to inferior treatments. We propose a principled method based on Bayesian hypothesis testing to mitigate this issue. Specifically, we introduce a point null hypothesis that postulates equal effectiveness of treatments. This induces shrinkage toward equal randomization probabilities, with the degree of shrinkage controlled by the prior probability of the null hypothesis. Equal randomization and Thompson sampling arise as special cases when the prior probability is set to one or zero, respectively. Simulated and real-world examples illustrate that the proposed method balances highly variable Thompson sampling with static equal randomization. A simulation study demonstrates that the method can mitigate issues with ordinary Thompson sampling and has comparable statistical properties to Thompson sampling with common ad hoc modifications such as power transformation and probability capping. We implement the method in the open-source R package brar, enabling experimenters to easily perform point null Bayesian RAR and support more effective randomization of patients.
Keywords: Adaptive trials, A/B testing, Bayes factor, Bayesian model averaging, sharp null hypothesis, spike-and-slab prior
@@ 1 Introduction
Response-adaptive randomization (RAR) methods, sometimes also called outcome-adaptive randomization methods, randomly allocate experimental units (e.g., patients or animals) to treatments in a manner that is informed by accumulating data (Thall and Wathen,, 2007; Berry et al.,, 2010; Grieve,, 2016; Robertson et al.,, 2023). A popular approach is Thompson sampling (Thompson,, 1933), which randomizes participants proportionally to the Bayesian posterior probability that each treatment is the most effective. Such RAR methods are attractive because they naturally balance gathering information on treatment effectiveness and assigning subjects to effective treatments.
Despite its benefits, there are also various challenges with Thompson sampling. For example, the method can exhibit high variability, particularly in the early stages of a study when posterior uncertainty is high (Thall et al.,, 2015). This can lead to erratic allocation probabilities, ethical concerns (e.g., exposing participants to inferior treatments with high probability), and inferential challenges (e.g., reduction of statistical power or biased effect estimates). Consequently, there has been substantial interest in modifying Thompson sampling to make it more reliable.
@@
For example, Thall and Wathen, (2007) propose to use the randomization probability $\pi=p^{c}/\{p^{c}+(1-p)^{c}\}$ where $p$ is the posterior probability that the experimental treatment is more effective than the control treatment, and $c$ is an additional parameter that controls the variability of the method. Setting $c=1$ produces Thompson sampling, while $c<1$ reduces variability with $c=1/2$ being often recommended. Another approach to reduce extreme randomization probabilities is to cap them, for instance, setting them to 10% or 90% if a method assigns more extreme probabilities (Thall and Wathen,, 2007; Lee and Lee,, 2021). Finally, RAR methods are often combined with “burn-in” periods at the start of the study, during which units are randomized with equal probabilities to mitigate high variability (Thall and Wathen,, 2007; Wathen and Thall,, 2017; Robertson et al.,, 2023).
@@
While such ad hoc modifications can address some of the limitations of RAR methods, they conflict with the principles of coherent Bayesian learning. For example, a transformed or capped posterior probability does no longer correspond to an actual posterior probability, and it cannot be used as a genuine prior for future data. This raises the question of whether it is possible to devise a RAR method with desirable properties, such as reduced variability compared to Thompson sampling, that is coherent with Bayesian principles, and if so, how it relates to these ad hoc modifications.
@@
In this paper, we propose a novel RAR method that reduces variability in a coherent Bayesian manner, which we term “point null Bayesian RAR”. The idea is to consider a point null hypothesis postulating that treatments are equally effective. This is equivalent to using a “spike-and-slab” prior (Raftery et al.,, 1997), also known as “lump-and-smear” prior, which is a mixture of a point mass at equal effectiveness and a probability density elsewhere (Spiegelhalter et al.,, 2004, Section 5.5.4). The prior probability of the point null hypothesis determines the mixture weight. As we will show, setting this prior probability to zero produces equal randomization, whereas setting it to one produces Thompson sampling. The proposed method thus interpolates between equal randomization and Thompson sampling in a coherent Bayesian way. As a by-product, posterior probabilities and Bayes factors are also obtained. These can be used to monitor evidence of the effectiveness of each treatment in a manner that is aligned with the randomization probabilities.
@@
In the following Section 2 we introduce the general idea of the method in more detail, followed by tailoring it to the setting of approximately normal effect estimates (Section 3) and binary outcomes (Section 4). In Section 5, we then illustrate the method on data from the ECMO trial (Bartlett et al.,, 1985), followed by evaluating its statistical properties in a simulation study (Section 6). The paper ends with concluding remarks on advantages, limitations, and opportunities for future research (Section 7). Appendix A illustrates our R package brar for performing point null Bayesian RAR. Appendix B provides further details on our simulation study.
2 Point null Bayesian RAR
We now explain the general idea of point null Bayesian RAR, without going into specifics such as data distribution or computation (these will follow in the subsequent sections). Throughout we will assume that we have observed data $y$ and we want to use these to randomize a future experimental unit. We start with the basic but important setting of one control and one treatment group, and extend it afterwards to multiple treatment groups.
2.1 Two group comparisons
In case there is a control group and only one treatment group, we consider the hypotheses:
@@ 2.1 Two group comparisons
$H_{+}\colon\text{Treatment is more effective than control}$
@@
How exactly these statements are translated into statistical hypotheses related to parameters depends on the type of data and model used, but often relates to an effect size parameter being less, equal, or greater than zero. There may also be situations where there is no control group but only two competing treatments. In this case, the method detailed here is still applicable but with the control group replaced by a reference group (the choice of the reference may be somewhat arbitrary).
The Bayesian posterior probability of a hypothesis $H_{i}\in\{H_{-},H_{0},H_{+}\}$ can then be calculated by
$\Pr(H_{i}\mid y)=\frac{p(y\mid H_{i})\Pr(H_{i})}{\sum\limits_{j\in\{-,0,+\}}p(y\mid H_{j})\Pr(H_{j})}=\left\{\sum_{j\in\{-,0,+\}}\mathrm{BF}_{ji}(y)\,\frac{\Pr(H_{j})}{\Pr(H_{i})}\right\}^{-1}$ (1)
where $\Pr(H_{i})$ is the prior probability of hypothesis $H_{i}$,
$p(y\mid H_{i})=\int_{\Theta}p(y\mid\theta)\,p(\theta\mid H_{i})\,\mathrm{d}\theta$
@@
is the marginal likelihood of the data $y$ under $H_{i}$ obtained from marginalizing the likelihood $p(y\mid\theta)$ with respect to the prior distribution $p(\theta\mid H_{i})$ assigned to the model parameters $\theta\in\Theta$ under $H_{i}$, and
$\mathrm{BF}_{ji}(y)=\frac{\Pr(H_{j}\mid y)}{\Pr(H_{i}\mid y)}\,\bigg/\,\frac{\Pr(H_{j})}{\Pr(H_{i})}=\frac{p(y\mid H_{j})}{p(y\mid H_{i})}$ (2)
@@
is the Bayes factor contrasting $H_{j}$ to $H_{i}$ (Jeffreys,, 1939; Good,, 1958). The Bayes factor (2) is the updating factor of the prior odds of $H_{j}$ to $H_{i}$ to the corresponding posterior odds (first equality), which is equivalent to the ratio of marginal likelihoods of the data under $H_{j}$ and $H_{i}$ (second equality). The posterior probabilities (1) can thus be computed from the marginal likelihoods of the data under each considered hypotheses along with their prior probabilities, or from the set of Bayes factors and prior hypothesis odds relative to some reference hypothesis (Kass and Raftery,, 1995).
@@
Regardless in which way they are computed, the question is how to translate posterior probabilities into randomization probabilities. We propose to randomize a future unit to the treatment group with probability
$\pi=\Pr(H_{+}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{2},$ (3)
while the probability to randomize to the control group is consequently
$1-\pi=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{2}.$ (4)
@@
Figure 1 shows such randomization probabilities for different combinations of posterior probabilities. We can see that the randomization probability $\pi$ shrinks towards 50% as the posterior probability of the null hypothesis $\Pr(H_{0}\mid y)$ increases. For example, if $\Pr(H_{+}\mid y)=0.2$, $\Pr(H_{-}\mid y)=0.3$, and $\Pr(H_{0}\mid y)=0.5$, as indicated by the black asterisk in Figure 1, the probability to randomize to the treatment group is $\pi=0.2+0.5/2=45\%$.
@@
Figure 1: Ternary plot where the color of each point depicts the treatment randomization probability $\pi$ for a particular combination of posterior probabilities of $H_{+}$ (bottom axis), $H_{-}$ (left axis), and $H_{0}$ (right axis). For example, the black asterisk denotes the combination where $\Pr(H_{+}\mid y)=0.2$, $\Pr(H_{-}\mid y)=0.3$, and $\Pr(H_{0}\mid y)=0.5$ with corresponding randomization probability $\pi=45\%$.
@@
This scheme can be motivated as Bayesian hypothesis averaged randomization probability where the hypothesis-specific treatment randomization probabilities are $0\%$, $50\%$, and $100\%$, under $H_{-}$, $H_{0}$, and $H_{+}$, respectively. That is, if we would know that $H_{+}$ or $H_{-}$ is true, we should assign the next patient to the treatment or control group, respectively, with probability one to maximize utility (patient benefit). On the other hand, if we knew that $H_{0}$ is true, then randomizing the assignment with $\pi=50\%$ seems sensible.
@@
Interestingly, the randomization scheme (3) reduces to Thompson sampling when the prior probability of $H_{0}$ is set to zero since then $\Pr(H_{0}\mid y)=0$ regardless of the data, and consequently $\pi=\Pr(H_{+}\mid y)$. However, the scheme induces shrinkage toward randomization probabilities of $\pi=50\%$ otherwise. In the most extreme case when $\Pr(H_{0})=1$, equal randomization is obtained as then $\Pr(H_{0}\mid y)=1$ regardless of the data, and consequently $\pi=50\%$. The scheme thus interpolates between Thompson sampling and equal randomization in a coherent Bayesian way.
@@ 2.2 More than two groups
$\vdots$
[…]
Posterior probabilities of each hypothesis can be computed from the marginal likelihoods and prior hypotheses probabilities with the summation in (1) extended to encompass all hypotheses (i.e., summing over $j\in\{-,0,+1,\dots,+K\}$). Similarly, they can be translated into randomization probabilities
$\pi_{i}=\Pr(H_{+i}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+1}$ (5)
with corresponding control randomization probability
$1-\sum_{i=1}^{K}\pi_{i}=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+1},$ (6)
which reduce to the randomization probabilities (3) and (4) for $K=1$.
@@
Also in the multi-treatment case, the randomization probabilities are shrunken towards equal randomization $\pi=1/(K+1)$ by introducing the null hypothesis $H_{0}$. Similarly, Thompson sampling and equal randomization are obtained by setting $\Pr(H_{0})=0$ and $\Pr(H_{0})=1$, respectively, since then the posteriors $\Pr(H_{0}\mid y)=0$ and $\Pr(H_{0}\mid y)=1$ are obtained for any observed data $y$.
@@
In this paper, we focus on schemes (5) and (6), which induce shrinkage toward equal randomization. However, it is important to note that there are alternatives to this approach. For example, it may be desired to shrink to different “baseline” randomization probabilities than equal randomization probabilities. This can be achieved by modifying the multiplicative factor of $\Pr(H_{0}\mid y)$ in (5) from $1/(K+1)$ to the desired baseline randomization probability. For example, if the goal is to minimize the standard errors of the treatment effect estimates by using $\sqrt{K}:1:\dots:1$ square-root allocation of the control to treatments (Dunnett,, 1955), we may use
$\pi_{i}=\Pr(H_{+i}\mid y)+\Pr(H_{0}\mid y)\times\frac{1}{K+\sqrt{K}}$
which leads to the control randomization probability
$1-\sum_{i=1}^{K}\pi_{i}=\Pr(H_{-}\mid y)+\Pr(H_{0}\mid y)\times\frac{\sqrt{K}}{K+\sqrt{K}}$
These correctly shrink RAR probabilitites towards the Dunnett-type randomization probabilities, $\pi_{i}=1/(K+\sqrt{K})$ for treatment $i=1,\dots,K$, and $1-\sum_{i=1}^{K}\pi_{i}=\sqrt{K}/(K+\sqrt{K})$ for control.
3 Point null Bayesian RAR under approximate normality
We will expand on point null Bayesian RAR in the setting where the data are summarized by an asymptotically normally distributed effect estimate. This does not mean that the raw data (e.g., a vector of outcomes and a matrix of covariates), from which the estimate is computed, need to be normally distributed. For instance, the regression coefficients from generalized linear models estimated using maximum likelihood satisfy asymptotic normality even if the data are themselves not normally distributed. This is useful because it allows us to efficiently compute posterior and randomization probabilities without the need for simulation. Although this framework is widely applicable, improvements may be possible by considering the exact distribution of the data. This will be detailed for binary outcomes in Section 4.
3.1 Two group comparisons

### v2

# Stabilizing Thompson Sampling with Null Hypothesis Bayesian Response-Adaptive Randomization

@@ Abstract
Response-adaptive randomization (RAR) methods can be used to adapt randomization probabilities based on accumulating data, aiming to increase the probability of allocating patients to effective treatments. A popular RAR method is Thompson sampling, which randomizes patients proportionally to the Bayesian posterior probability that each treatment is the most effective. However, its high variability can also increase the risk of assigning patients to inferior treatments and lead to inferential problems such as confidence interval undercoverage. We propose a principled method based on Bayesian hypothesis testing to address these issues: We introduce a null hypothesis postulating equal effectiveness of treatments. Bayesian model averaging then induces shrinkage toward equal randomization probabilities, with the degree of shrinkage controlled by the prior probability of the null hypothesis. Equal randomization and Thompson sampling arise as special cases when the prior probability is set to one or zero, respectively. Simulated and real-world examples illustrate that the method balances highly variable Thompson sampling with static equal randomization. A simulation study demonstrates that the method can mitigate issues with Thompson sampling and has comparable statistical properties to Thompson sampling with common ad hoc modifications such as power transformation and probability capping. We implement the method in the free and open-source R package brar, enabling experimenters to easily perform null hypothesis Bayesian RAR and support more effective randomization of patients.
Keywords: Adaptive trials, Bayes factor, Bayesian model averaging, outcome-adaptive randomization, spike-and-slab prior
@@ 1 Introduction
Response-adaptive randomization (RAR) methods randomly allocate patients to treatments based on accumulating data (Hu and Rosenberger,, 2006; Thall and Wathen,, 2007; Berry et al.,, 2010; Robertson et al.,, 2023). A popular approach is Thompson sampling (Thompson,, 1933), which randomizes patients proportionally to the Bayesian posterior probability that each treatment is the most effective. Such RAR methods are attractive as they balance gathering information on treatment effectiveness and assigning subjects to effective treatments.
Despite its benefits, Thompson sampling can also introduce inferential problems such as inflated type I error rates or bias in effect estimation. Furthermore, although Thompson sampling allocates patients to the more effective treatment on average, it may also increase the probability of allocating patients to inferior treatments compared to equal randomization, particularly when treatment effects are small. This can undermine the method’s intended objective and raise ethical concerns (Thall et al.,, 2015; Robertson et al.,, 2023). Consequently, there has been interest in modifying Thompson sampling to address these issues.
@@
One popular modification uses the randomization probability $\pi=p^{c}/\{p^{c}+(1-p)^{c}\}$, where $p$ is the posterior probability that the experimental treatment is more effective than the control, and $c$ is a tuning parameter. Setting $c=1$ produces Thompson sampling while $c<1$ reduces variability. Another approach is to cap randomization probabilities, for instance, setting them to 10% or 90% if the method assigns more extreme probabilities. Finally, “burn-in” periods at the start of the study are commonly used, during which patients are randomized with equal probabilities to mitigate high variability (Wathen and Thall,, 2017).
@@
Several simulation studies have shown that appropriately modified Thompson sampling can overcome the limitations of the ordinary formulation (e.g., Du et al.,, 2017; Viele et al.,, 2019, 2020; Lee and Lee,, 2021; Tang et al.,, 2025). However, such ad hoc modifications conflict with principles of coherent Bayesian learning. For example, a capped posterior no longer corresponds to an actual posterior and cannot serve as a genuine prior for future data. This raises the question of whether a principled Bayesian RAR method can mitigate Thompson sampling’s variability, and if so, how it relates to ad hoc modifications.
Principled Bayesian alternatives to Thompson sampling have also been proposed outside clinical trials, particularly in the multi-armed bandit literature. For instance, RAR based on the Gittins index prioritizes treatments with the greatest expected reward (Gittins et al.,, 2011), or the Bayesian upper confidence bound method allocates treatments based on posterior quantiles and offers theoretical performance guarantees (Kaufmann et al.,, 2012). These methods are primarily used in online decision problems but rarely in clinical trials, where additional constraints such as type I error rate control are central (Villar et al.,, 2015).
@@
In this paper, we propose a novel RAR method that reduces variability in a coherent Bayesian manner, which we term “null hypothesis Bayesian RAR”. The idea is to introduce a null hypothesis postulating that treatments are equally effective. This is often equivalent to using a “spike-and-slab” prior for the treatment effect parameter, which is a mixture of a point mass at equal effectiveness and a probability density elsewhere (Spiegelhalter et al.,, 2004, Section 5.5.4). The prior probability of the null hypothesis determines the mixture weight. Setting it to zero produces equal randomization, whereas setting it to one produces Thompson sampling. The method thus interpolates between equal randomization and Thompson sampling in a coherent Bayesian way. As a by-product, posterior probabilities and Bayes factors are also obtained. These can be used to monitor evidence of treatment effectiveness in a manner that is aligned with the randomization probabilities.
@@
In the following Section 2 we introduce the general idea of the method, followed by tailoring it to normal (Section 3) and binary outcomes (Section 4). Section 5 illustrates the method on data from the ECMO trial, followed by evaluating its properties in a simulation study (Section 6). Concluding remarks are given in Section 7. The supplement illustrates our R package brar, and provides further details on theory and the simulation study.
2 Null hypothesis Bayesian RAR
Assume that we have observed data $y$ and want to use these to randomize a future patient. Suppose there is one control and one treatment group, and consider the hypotheses:
@@ 2 Null hypothesis Bayesian RAR
$H_{+}\colon\text{Treatment is more effective than control.}$
@@
How exactly these are translated into statistical hypotheses about parameters depends on the data model, but often relates to an effect size parameter being less, equal, or greater than zero. If there is no control group, the method can still be used with the control replaced by a reference group. This hypothesis setup is similar to Cai et al., (2013) but with the addition of $H_{0}$, which introduces a different behavior of the resulting randomization probabilities.
The Bayesian posterior probability of hypothesis $H_{i}\in\{H_{-},H_{0},H_{+}\}$ is then
$\Pr(H_{i}\,|\,y)=\frac{p(y\,|\,H_{i})\Pr(H_{i})}{\sum\limits_{j\in\{-,0,+\}}p(y\,|\,H_{j})\Pr(H_{j})}=1\,\bigg/\,\bigg\{\sum_{j\in\{-,0,+\}}\mathrm{BF}_{ji}(y)\,\frac{\Pr(H_{j})}{\Pr(H_{i})}\bigg\}$ (1)
@@
where $\Pr(H_{i})$ is the prior probability of $H_{i}$ (with $\sum_{i\in\{-,0,+\}}\Pr(H_{i})=1$), $p(y\,|\,H_{i})=\int_{\Theta}p(y\mid\theta)\,p(\theta\,|\,H_{i})\,\mathrm{d}\theta$ is the marginal likelihood of the data $y$ under $H_{i}$ obtained from integrating the likelihood $p(y\,|\,\theta)$ with respect to the prior distribution $p(\theta\,|\,H_{i})$ of the model parameters $\theta\in\Theta$ under $H_{i}$, and
$\mathrm{BF}_{ji}(y)=\frac{\Pr(H_{j}\,|\,y)}{\Pr(H_{i}\,|\,y)}\,\bigg/\,\frac{\Pr(H_{j})}{\Pr(H_{i})}=\frac{p(y\,|\,H_{j})}{p(y\,|\,H_{i})}$ (2)
@@
is the Bayes factor contrasting $H_{j}$ to $H_{i}$ (Kass and Raftery,, 1995). The Bayes factor (2) is the updating factor of the prior odds of $H_{j}$ to $H_{i}$ to the corresponding posterior odds (first equality), which is equivalent to the ratio of marginal likelihoods of the data under $H_{j}$ and $H_{i}$ (second equality). The posterior probabilities (1) can thus be computed from the marginal likelihoods of the data under each considered hypotheses along with their prior probabilities, or from the Bayes factors and prior odds relative to some reference hypothesis.
@@
Regardless in which way they are computed, the question is how to translate posterior probabilities into randomization probabilities. We propose to randomize a future patient to the treatment group with probability
$\pi=\Pr(H_{+}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,2,$ (3)
@@
while the probability to randomize to the control group is consequently $1-\pi=\Pr(H_{-}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,2$. Figure 1 shows $\pi$ for different combinations of posterior probabilities. We can see that $\pi$ shrinks towards 50% as the posterior probability of the null hypothesis $\Pr(H_{0}\,|\,y)$ increases. For example, if $\Pr(H_{+}\,|\,y)=0.2$, $\Pr(H_{-}\,|\,y)=0.3$, and $\Pr(H_{0}\,|\,y)=0.5$, as indicated by the black asterisk, the randomization probability is $\pi=0.2+0.5/2=45\%$.
@@
Figure 1: Ternary plot where the color of each point depicts the treatment randomization probability $\pi$ for a particular combination of posterior probabilities of $H_{+}$ (bottom axis), $H_{-}$ (left axis), and $H_{0}$ (right axis). For example, the black asterisk denotes the combination where $\Pr(H_{+}\,|\,y)=0.2$, $\Pr(H_{-}\,|\,y)=0.3$, and $\Pr(H_{0}\,|\,y)=0.5$ with corresponding randomization probability $\pi=45\%$.
@@
This scheme can be motivated as Bayesian hypothesis averaged randomization probability where the hypothesis-specific treatment randomization probabilities are $0\%$, $50\%$, and $100\%$, under $H_{-}$, $H_{0}$, and $H_{+}$, respectively. That is, if we knew that $H_{+}$ or $H_{-}$ is true, we should assign the next patient to the treatment or control group, respectively, to maximize patient benefit. In contrast, if we knew that $H_{0}$ is true, randomizing with $\pi=50\%$ seems sensible.
@@
The randomization scheme (3) reduces to Thompson sampling when the prior probability of $H_{0}$ is set to zero since then $\Pr(H_{0}\,|\,y)=0$ regardless of the data, and consequently $\pi=\Pr(H_{+}\,|\,y)$. However, the scheme induces shrinkage toward $\pi=50\%$ otherwise. In the most extreme case when $\Pr(H_{0})=1$, equal randomization is obtained as then $\Pr(H_{0}\,|\,y)=1$ regardless of the data, and consequently $\pi=50\%$. The scheme thus interpolates between Thompson sampling and equal randomization in a coherent Bayesian way.
@@ 2.1 More than two groups
for $i=1,\dots,K,$ which span all effect ordering possibilities. Posterior hypothesis probabilities can still be computed from (1) but summing over all hypotheses (i.e., $j\in\{-,0,+1,\dots,+K\}$). These can then be translated into treatment randomization probabilities
$\pi_{i}=\Pr(H_{+i}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,(K+1)$ (4)
and control randomization probability $1-\sum_{i=1}^{K}\pi_{i}=\Pr(H_{-}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,(K+1)$.
@@
Also in the multi-treatment case, the randomization probabilities shrink towards equal randomization $\pi_{i}=1/(K+1)$ by introducing the null hypothesis $H_{0}$. Thompson sampling and equal randomization are again obtained by setting $\Pr(H_{0})=0$ and $\Pr(H_{0})=1$, respectively, as this leads to $\Pr(H_{0}\,|\,y)=0$ and $\Pr(H_{0}\,|\,y)=1$ for any data $y$.
@@
In this paper, we focus on scheme (4), which induces shrinkage toward equal randomization. However, it is important to note that there are alternatives. It may be desired to shrink to different “baseline” randomization probabilities. This can be achieved by modifying the factor of $\Pr(H_{0}\,|\,y)$ in (4) from $1/(K+1)$ to the desired baseline randomization probability. For example, the aim may be to test all $K$ pairwise differences between the treatments and the common control using Dunnett’s test. A square root allocation rule (with $\sqrt{K}$ more patients in the control group than in each treatment group) is commonly used to maximize the power of the test (Kieser,, 2020, p.156). We may therefore use $\pi_{i}=\Pr(H_{+i}\,|\,y)+\Pr(H_{0}\,|\,y)\,/\,(K+\sqrt{K})$. These correctly shrink RAR probabilitites towards Dunnett-type randomization probabilities, $\pi_{i}=1/(K+\sqrt{K})$ for treatment $i=1,\dots,K$, and $1-\sum_{i=1}^{K}\pi_{i}=\sqrt{K}/(K+\sqrt{K})$ for control.
2.2 Asymptotic randomization probabilities
It is of interest to understand how RAR probabilities behave as data accumulate. Under ordinary non-RAR based sampling, the posterior distribution is “consistent”, meaning that as more data are generated under a hypothesis the corresponding posterior hypothesis probability tends to one (Dawid,, 2011). In Section 3, we provide empirical support through simulation that posterior consistency also holds for the RAR setting. Thus if a treatment is superior and $\Pr(H_{0})<1$, the corresponding posterior and randomization probability tend to one. Conversely, if all treatments are equally effective and $\Pr(H_{0})>0$, the posterior probability of $H_{0}$ will approach one, and the randomization probabilities will hence tend to the specified baseline randomization probabilities. This is an important difference to Thompson sampling, whose randomization probabilities do not converge to the baseline randomization probability under $H_{0}$ but remain random (Deliu and Villar,, 2025).
3 Null hypothesis Bayesian RAR under normality